SEARCH FOR THE LOW-LYING ELECTRONICALLY EXCITED STATES OF UO_{2}

\underline{I. O. ANTONOV}, M. C. HEAVEN, Department of Chemistry, Emory University, Atlanta, GA 30322.

Low-lying electronically excited states of UO_{2} were studied by laser excitation and dispersed fluorescence spectroscopy. Excitation of UO_{2} in 31300-31800 cm^{−1} and 27100 cm^{−1} region produced highly congested emission spectrum. A search for the low-lying absorption bands was performed in the region of 10000-15000 cm^{−1} using two-color REMPI spectroscopy. Several band systems were found near 11500, 12300, 13100 and 13700 cm^{−1}.